### v1

# ClearVision: Leveraging CycleGAN and SigLIP-2 for Robust All-Weather Classification in Traffic Camera Imagery

Abstract
Accurate weather classification from low-quality traffic camera imagery remains a challenging task, particularly under adverse night-time conditions. In this study, we propose a scalable framework that combines generative domain adaptation with efficient contrastive learning to enhance classification performance. Using CycleGAN-based domain translation, we improve the quality of night-time images, enabling better feature extraction by downstream models. While the baseline EVA-02 model employing CLIP-based contrastive loss achieves an overall accuracy of 96.55%, it shows a significant performance gap between day-time (97.21%) and night-time conditions (63.40%). Replacing CLIP with the more lightweight SigLIP-2 (Sigmoid contrastive loss) achieves competitive overall accuracy of 94.00%, with significant improvements in night-time performance (85.90% accuracy). The combination of Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive training achieves the best night-time accuracy (85.90%) across all models tested, while EVA-02 with CycleGAN maintains the highest overall accuracy (97.01%) and per-class accuracies. Our findings demonstrate the potential of combining domain adaptation and efficient contrastive learning to build practical, resource-efficient weather classification systems for intelligent transportation infrastructure.
I INTRODUCTION
Figure 1: SigLIP-2 + CycleGAN based Architecture
Adverse weather contributes to around 1.2 million traffic accidents annually in the U.S.[1], making timely, localized weather detection crucial for safety[2, 3]. Traditional methods, like satellite imagery and meteorological stations [4, 5], lack the spatial resolution needed for real-time road-specific insights, especially during precipitation events that impair visibility and traction [6]. To address this limitation, computer vision uses roadside traffic cameras to provide a scalable, cost-effective solution for real-time weather detection. However, CNN-based models used in prior approaches suffer from generalizability issues [7], heavily relying on large, diverse datasets [8, 9].
In response to these challenges, Vision Transformers (ViTs) have emerged as a robust alternative to traditional CNN-based models. By leveraging self-attention mechanisms, ViTs can model local and global dependencies within images [10], offering superior pattern recognition capabilities. Abdelraouf et al. [11] proposed a ViT architecture with a self-spatial attention module for rain and road-surface classification, achieving strong binary classification results. More recently, Chen et al. [12] introduced MASK-CT, a hybrid model combining masked convolutional networks with transformers to enhance generalization. Earlier approaches using conventional architectures like AlexNet and ResNet-18 [13, 3] achieved reasonable accuracy but struggled with broader weather pattern recognition and generalization across diverse camera views [14, 15]. While ViT-based models offer strong performance, they often require substantial computational resources and high-resolution training data, an issue given that real-world traffic camera footage is typically low in quality, particularly under nighttime or adverse weather conditions.
In this paper, we propose a novel framework that addresses both the computational burden and poor nighttime performance limitations of existing approaches. Our solution centers on Sigmoid Loss for Language-Image Pre-training (SigLIP-2-2) [16], which uses a pairwise sigmoid loss instead of the conventional cross-entropy across batch samples. This significantly reduces computational requirements while maintaining strong performance. We integrate SigLIP-2-2 with a CycleGAN-based domain adaptation technique within a contrastive learning framework, creating a comprehensive solution specifically designed for weather classification from traffic camera imagery, as illustrated in Fig. 1.
The core innovation of our approach lies in the synergistic combination of these components. SigLIP-2-2 provides efficient visual-textual representation learning, CycleGAN enhances nighttime frames through domain translation, and contrastive learning further refines the embedding space to maximize discrimination between weather conditions. Using a roadside camera dataset from the Iowa Department of Transportation, we demonstrate that our Vision-SigLIP-2-2 + Text-SigLIP-2-2 + CycleGAN + Contrastive Learning framework achieves the best nighttime performance (85.90% accuracy) across all tested models while maintaining strong overall accuracy (94.00%).
We benchmark our proposed framework against established models, including EVA-02 with CLIP and a standard Vision Transformer (vit-base-patch16-224-in21k) for a comprehensive evaluation. While EVA-02 with CLIP achieves slightly higher overall accuracy (97.01%), it comes at significantly higher computational cost. EVA-02 integrates CLIP with Transformer-based architectures [17] to achieve state-of-the-art performance, but relies on computationally intensive batch-wise softmax operations and the heavy memory footprint of InfoNCE loss during training [18], making it less suitable for resource-constrained deployment scenarios.
Key Contributions: This paper makes several notable contributions to the field of weather detection using traffic camera imagery:
  1. 1.
    We propose a lightweight alternative to conventional CLIP-based models by incorporating SigLIP-2-2 for efficient language-image pretraining, significantly reducing computational overhead while maintaining competitive accuracy.
  2. 2.
    We introduce an innovative CycleGAN-based domain adaptation technique specifically designed for traffic camera imagery that enhances nighttime frames, demonstrating substantial performance improvements under low-light conditions (up to 19.05% accuracy gain).
  3. 3.
    We develop and evaluate a novel contrastive learning framework that further improves the performance of SigLIP-2-based models, particularly for challenging weather conditions and low-light scenarios.
  4. 4.
    We present comprehensive benchmarking of various model configurations on a real-world traffic camera dataset, providing valuable insights into the trade-offs between computational efficiency and accuracy for weather classification tasks.
  5. 5.
    We demonstrate that while EVA-02 with CLIP achieves the highest accuracy (97.01%), our SigLIP-2-based approaches offer a more favorable balance between performance (94.00% accuracy) and computational efficiency for practical deployments.
  6. 6.
    We achieve significant computational efficiency gains with our SigLIP-2-based approach, reducing training time by 89% and inference time by 83% compared to EVA-02, making our solution ideal for resource-constrained real-world applications.
Paper Organization: The remainder of this paper is organized as follows: Section II-A1 establishes the mathematical foundations of our approach and details the loss functions used for training. Sections II-B1 and II-B3 describe the EVA-02 and CycleGAN architectures, respectively. Section III presents our dataset and training protocols. Section IV comprehensively analyzes experimental results, including model performance comparisons, CycleGAN enhancement effects, and qualitative analysis. Finally, we conclude with a discussion of limitations and directions for future work in Section V.
II Methods
In this section, we present our framework integrating SigLIP-2 with CycleGAN for robust weather classification from traffic camera imagery.
II-A Mathematical Formulation
II-A1 Task and Pipeline
[…]
While enhancing night-time scenes:
  1. (i)
    Initial classification: pretrained SigLIP-2 provides the first decision.
  2. (ii)
    Fine-tuning: SigLIP-2 and CycleGAN optimized on mis-classified samples.
  3. (iii)
    Enhancement: CycleGAN converts night images into day-like renderings.
  4. (iv)
    Re-classification: fine-tuned SigLIP-2 revisits enhanced images.
Contrastive learning constrains the embedding space to remain discriminative.
II-A2 Input and Output
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, $x_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$, $y_{i}\!\in\!\mathcal{Y}$, and an unpaired set $\mathcal{D}_{\mathrm{unpaired}}=\{x_{j}^{\mathrm{night}},\,x_{k}^{\mathrm{day}}\}$. The network predicts
$p(y_{i}\!\mid\!x_{i})\in[0,1]^{3},\quad\sum_{c=1}^{3}p(y_{i}\!=\!c\mid x_{i})=1.$
II-A3 Model Definitions
  • SigLIP-2 encoder: $f_{\bm{\theta}}\!:\mathbb{R}^{H\times W\times 3}\!\to\!\mathbb{R}^{768}$
  • Projection head: $h_{\bm{\phi}}\!:\mathbb{R}^{768}\!\to\!\mathbb{R}^{128}$
  • Classification head: $c_{\bm{\psi}}\!:\mathbb{R}^{768}\!\to\![0,1]^{3}$
  • CycleGAN generators: $G_{\bm{\alpha}}\!:X^{\mathrm{night}}\!\to\!Y^{\mathrm{day}}$,  $F_{\bm{\beta}}\!:Y^{\mathrm{day}}\!\to\!X^{\mathrm{night}}$
  • Discriminators: $D_{X}$, $D_{Y}$
II-A4 Architecture Details
SigLIP-2 Framework
Our proposed SigLIP-2 framework employs a dual-encoder architecture similar to CLIP but with important modifications for improved efficiency. The vision encoder processes images as $16\times 16$ pixel patches with 12 layers and 12 attention heads, producing 768-dimensional embeddings. A similar transformer design is used for the text encoder, processing weather condition descriptions.
The key innovation of SigLIP-2 compared to CLIP is its training objective. While CLIP uses InfoNCE loss requiring large batch sizes and significant GPU memory, SigLIP-2 employs a more efficient pairwise sigmoid-based contrastive approach, enabling equivalent performance with significantly reduced computational requirements—critical for practical deployment in transportation infrastructure.
CycleGAN Architecture
The CycleGAN component consists of two generator networks ($G_{X\rightarrow Y}$ and $G_{Y\rightarrow X}$) that learn mappings between night and day domains, and two discriminator networks ($D_{X}$ and $D_{Y}$) that distinguish between real and generated images. The framework maintains cycle-consistency to ensure that translating an image from one domain to another and back preserves the original content while modifying only domain-specific features.
II-A5 Loss Functions
Define the error set =$\mathcal{M}=\{x_{i}\mid\arg\!\max_{c}c_{\bm{\psi}}(f_{\bm{\theta}}(x_{i}))_{c}\neq y_{i}\}$, its CycleGAN outputs $\tilde{x}_{i}=G_{\bm{\alpha}}(x_{i})$, and $\mathcal{X}=\{x_{i}\}_{i=1}^{N}\cup\{\tilde{x}_{i}\}$.
(1) Classification loss
$\mathcal{L}_{\mathrm{cls}}$ $=$ $-\frac{1}{|\mathcal{X}|}\!\sum_{x\in\mathcal{X}}\sum_{c=1}^{3}\bigl{[}(1-\varepsilon)\mathbf{1}(y_{x}{=}c)+\tfrac{\varepsilon}{3}\bigr{]}$ (1)
×${}\times\log c_{\bm{\psi}}\!\bigl{(}f_{\bm{\theta}}(x)\bigr{)}_{c},$
where $\varepsilon=0.1$ (label smoothing).
(2) Contrastive loss
$\mathcal{L}_{\mathrm{con}}$ $=$ $\frac{1}{|\mathcal{P}|}\!\sum_{(i,j)\in\mathcal{P}}\Bigl{[}-\!\log\sigma\!\bigl{(}\tfrac{\bm{e}_{i}^{\!\top}\bm{e}_{j}}{\tau}\bigr{)}$ (2)
${}\;-\!\sum_{k:y_{k}\neq y_{i}}\log\!\Bigl{(}1-\sigma\!\bigl{(}\tfrac{\bm{e}_{i}^{\!\top}\bm{e}_{k}}{\tau}\bigr{)}\Bigr{)}\Bigr{]},$
where $\bm{e}_{i}=h_{\bm{\phi}}(f_{\bm{\theta}}(x_{i}))$, $\tau=0.1$, and $\mathcal{P}=\{(x_{i},x_{j})\mid y_{i}=y_{j}\}\cup\{(x_{i},\tilde{x}_{i})\mid x_{i}\in\mathcal{M}\}$.
(3) CycleGAN loss
$\mathcal{L}_{\mathrm{cycGAN}}=\mathcal{L}_{\mathrm{adv}}(G_{\bm{\alpha}},D_{Y})+\mathcal{L}_{\mathrm{adv}}(F_{\bm{\beta}},D_{X})$
$\quad+\,\lambda_{\mathrm{cyc}}\mathcal{L}_{\mathrm{cyc}}+\lambda_{\mathrm{id}}\mathcal{L}_{\mathrm{id}}+\lambda_{\mathrm{weather}}\mathcal{L}_{\mathrm{weather}},$ (3)
with $(\lambda_{\mathrm{cyc}},\lambda_{\mathrm{id}},\lambda_{\mathrm{weather}})=(10,5,1)$ and
$\mathcal{L}_{\mathrm{adv}}(G,D_{Y})=\mathbb{E}_{y\sim p_{Y}}[\log D_{Y}(y)]$ (4)
$\quad+\mathbb{E}_{x\sim p_{X}}[\log(1-D_{Y}(G(x)))],$
$\mathcal{L}_{\mathrm{cyc}}=\mathbb{E}_{x\sim p_{X}}\!\bigl{\|}F(G(x))-x\bigr{\|}_{1}$ (5)
$\quad+\mathbb{E}_{y\sim p_{Y}}\!\bigl{\|}G(F(y))-y\bigr{\|}_{1},$
$\mathcal{L}_{\mathrm{id}}=\mathbb{E}_{x\sim p_{X}}\!\bigl{\|}G(x)-x\bigr{\|}_{1}$ (6)
$\quad+\mathbb{E}_{y\sim p_{Y}}\!\bigl{\|}F(y)-y\bigr{\|}_{1},$
$\mathcal{L}_{\mathrm{weather}}=\frac{1}{|\mathcal{M}|}\sum_{x_{i}\in\mathcal{M}}\sum_{c=1}^{3}\mathbf{1}(y_{i}{=}c)$ (7)
$\quad\cdot\log c_{\bm{\psi}}\!\bigl{(}f_{\bm{\theta}}(G_{\bm{\alpha}}(x_{i}))\bigr{)}_{c}.$ (8)
(4) Total loss
$\mathcal{L}_{\mathrm{total}}=\lambda_{\mathrm{con}}\mathcal{L}_{\mathrm{con}}+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}},\quad(\lambda_{\mathrm{con}},\lambda_{\mathrm{cls}})=(1,0.5).$
II-A6 Integrated Framework
The core contribution of our work is the integration of these components into a unified framework. As illustrated in Fig. 1, our system operates through the following pipeline:
  1. 1.
    Domain Translation: Night-time images undergo enhancement via CycleGAN to their day-time equivalents while preserving weather-specific features.
  2. 2.
    Feature Extraction: Both original and transformed images are processed through SigLIP-2 to extract rich visual features.
  3. 3.
    Contrastive Alignment: The contrastive learning framework ensures consistency between original and transformed representations, improving robustness to lighting variations.
  4. 4.
    Classification: The final classification layer produces probability distributions over three weather classes.
Our novel weather-preserving loss ($\mathcal{L}_{\mathrm{weather}}$) ensures the CycleGAN transformation maintains critical weather-related visual cues, addressing both computational efficiency constraints and challenging low-light conditions of traffic camera imagery.
II-B Implementation Details
II-B1 EVA-02 Transformer
EVA-02 builds upon the Vision Transformer architecture with multi-head self-attention for capturing spatial dependencies and position-wise feedforward networks for feature transformation. It incorporates Swish Gated Linear Unit activation [19], sub-Layer Normalization, and 2D Rotary Position Embedding. Unlike CNNs, EVA-02 processes images as sequences of patches, enabling global relationship learning without convolutional biases [17]. It achieves parameter efficiency through optimized attention mechanisms and reduced hidden layer dimensionality.
II-B2 Proposed SigLIP-2 Framework
Our framework centers on SigLIP-2, which replaces CLIP’s computationally expensive cross-entropy loss with efficient pairwise sigmoid loss. We enhance this with contrastive learning that maximizes discriminative power by:
1. Applying strong/weak augmentations to create multiple views of images 2. Ensuring embeddings of the same weather condition cluster together 3. Enforcing separation between different weather conditions
This approach is particularly effective for distinguishing similar weather appearances and improving night-time image classification after CycleGAN enhancement.
II-B3 Cycle Generative Adversarial Networks
CycleGANs learn bidirectional mappings between day and night domains without requiring paired examples [20]. The framework uses two generators ($G_{X\rightarrow Y}$ and $G_{Y\rightarrow X}$) and two discriminators ($D_{X}$ and $D_{Y}$), with cycle-consistency constraints ensuring content preservation while altering only domain-specific characteristics.
The core innovation is the cycle-consistency loss which enforces that an image translated from domain X to Y and back should match the original:
$\mathcal{L}_{\text{cyc}}(G_{X\rightarrow Y},G_{Y\rightarrow X})=\mathbb{E}_{x\sim p_{\text{data}}(X)}\left[\|G_{Y\rightarrow X}(G_{X\rightarrow Y}(x))-x\|_{1}\right]$ (9)
This is critical for our application, as the model must avoid introducing artificial precipitation artifacts during domain conversion. Traditional supervised approaches would require aligned day-night pairs under identical weather conditions—impractical to collect. A conceptual illustration appears in Fig. 2.
Figure 2: CycleGAN architecture for night-to-day image translation: The framework employs two generator networks ($G$ and $F$) that learn mappings between night and day domains, and two discriminator networks ($D_{X}$ and $D_{Y}$) that distinguish between real and generated images in each domain. The cycle-consistency constraint ensures that translating an image from one domain to another and back preserves the original content while modifying only domain-specific features.
II-B4 Integrated SigLIP-2 + CycleGAN Framework
Our integrated system creates a synergistic pipeline where CycleGAN pre-processes challenging night-time data, while SigLIP-2 with its efficient contrastive approach enables precise classification with reduced computational demands. The weather-preserving loss further ensures that translations maintain critical weather cues, creating a balanced solution for both computational efficiency and challenging visual conditions in transportation monitoring systems.
III Dataset
We used traffic camera imagery from CCTV installations in Ames, Iowa (Iowa DoT) spanning three weather conditions: No Precipitation, Rain, and Snow. Images were standardized to 224×224 pixels and filtered following [21, 14]. Models were adapted to address challenges typical in traffic imagery, including variable lighting and weather-induced noise [15].
Training protocols:
  • EVA-02 (with CLIP): Fine-tuned on 11,178 images (87.4/8.4/4.2% train/val/test split) across three weather classes using AdamW optimizer and cosine schedule [17].
  • SigLIP-2: Used identical dataset split and training parameters as the CLIP variant, representing our primary contribution for efficient contrastive learning.
  • Vision Transformer: Trained on a reduced subset (2,391 images) with 60-20-20 split due to computational constraints [22].
  • CycleGAN: Our domain adaptation component, trained on 2,204 unpaired day-night image sets across all weather conditions. Training used Adam optimizer (learning rate 0.0002, linear decay) with identity mapping loss weighted at 0.5 times the adversarial loss to preserve critical weather features [20].
(a) Generator Loss
(b) Discriminator Loss
(c) Combined Loss
Figure 3: CycleGAN training losses over 100 epochs showing stable convergence patterns typical of well-trained GAN models [23].
Training batches included adjacent frames where possible to enhance temporal awareness in weather pattern recognition.
IV Results and Discussion
The test set comprised a total of 4,564 images, with 1,452 in the No Precipitation class, 1,590 in the Rain class, and 1,522 in the Snow class. The performance results for all model configurations are presented in detail in Table I (overall, day/night performance) and Table II (per-class performance).
IV-A Model Performance Analysis
The baseline EVA-02 model with CLIP demonstrates strong overall performance with 96.55% accuracy, 96.80% precision, and 96.65% $F_{1}$ score. However, it shows a significant performance gap between daytime (97.21% accuracy) and nighttime conditions (63.40% accuracy), highlighting the challenge of low-light imagery classification. When enhanced with CycleGAN preprocessing, EVA-02 shows substantial improvement in night-time performance (82.45% accuracy) while maintaining strong day-time results (97.45% accuracy), resulting in the best overall performance among all models with 97.01% accuracy.
The Vision-SigLIP-2 + Text-SigLIP-2 models demonstrate promising results, starting with 87.00% overall accuracy in the base configuration. Adding CycleGAN improves this to 91.00%, with substantial gains in night-time performance (from 67.00% to 81.00% accuracy). Most notably, the combination of Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive achieves the best night-time performance across all models (85.90% accuracy) and competitive overall performance at 94.00% accuracy.
In the per-class analysis (Table II), while EVA-02 models achieve the highest per-class accuracies (No Precipitation: 98.10%, Rain: 97.35%, Snow: 95.60% with CycleGAN), the Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive configuration still demonstrates strong performance (No Precipitation: 96.80%, Rain: 93.50%, Snow: 92.10%), particularly considering its significantly lower computational requirements.
IV-B Vision Transformer Performance Analysis
The Vision Transformer models perform notably worse than both EVA-02 and SigLIP-2 variants, with overall accuracy of only 55.81% for the base model and 54.20% with CycleGAN. These models show particularly weak performance in the No Precipitation class (19.15% and 4.67% accuracy, respectively) but relatively better performance in rain conditions.
This significant performance degradation can be attributed to three main factors: (1) the low-quality, noisy nature of traffic camera images disrupts the self-attention mechanisms in Vision Transformers, (2) the lack of extensive pretraining compared to EVA-02 and SigLIP-2 models limits their generalization capability, and (3) the absence of inductive biases in Vision Transformers makes them more sensitive to the quality of training data. Without clear spatial structure and fine-grained features, the models struggle to correctly weigh the importance of image regions, leading to frequent misclassifications.
TABLE I: Accuracy, precision, and $F_{1}$ score (%) for all evaluated models under day-time and night-time lighting. The best value in each column is shown in bold.
Accuracy Precision $F_{1}$ score
Model Overall Day Night Overall Day Night Overall Day Night
EVA02 (baseline) 96.55 97.21 63.40 96.80 97.45 62.10 96.65 97.33 62.70
EVA02 + CycleGAN 97.01 97.45 82.45 97.25 97.66 82.10 97.10 97.55 82.26
Vision Transformer 55.81 59.05 52.50 59.79 59.78 63.54 59.13 60.16 59.96
Vision Transformer + CycleGAN 56.46 58.57 54.20 57.68 59.24 61.42 58.42 59.68 59.96
Vision-SigLIP-2 + text-SigLIP-2 87.00 89.40 67.00 87.50 90.00 66.00 87.20 89.70 66.50
Vision-SigLIP-2 + text-SigLIP-2 + CycleGAN 91.00 92.50 81.00 91.30 92.80 80.50 91.10 92.60 80.75
Vision-SigLIP-2 + text-SigLIP-2 + CycleGAN + Contrastive 90.35 94.80 85.90 90.25 95.00 85.50 90.30 94.90 85.70
TABLE II: Per-class accuracy, precision, and $F_{1}$ score (%) for every model. Best value in each column is in bold.
Accuracy (%) Precision (%) $F_{1}$ score (%)
Model Overall No Precip. Rain Snow Overall No Precip. Rain Snow Overall No Precip. Rain Snow
EVA02 (baseline) 96.55 97.40 96.20 96.10 96.80 97.98 96.63 94.88 96.65 97.39 96.15 96.16
EVA02 + CycleGAN 97.01 98.10 97.35 95.60 97.25 98.40 97.70 95.00 97.10 98.25 97.52 95.30
Vision Transformer 55.81 19.15 92.42 64.01 59.79 51.24 46.62 81.50 59.13 27.88 61.96 71.71
Vision Transformer + CycleGAN 54.20 4.67 88.39 82.70 61.42 67.35 43.01 73.91 48.22 8.73 57.87 78.05
Vision-SigLIP-2 + Text-SigLIP-2 87.00 91.50 85.60 83.90 87.50 92.10 86.10 84.40 87.20 91.80 85.85 84.15
Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN 91.00 94.60 90.10 88.30 91.30 94.90 90.50 88.70 91.10 94.75 90.30 88.50
Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive 94.00 96.80 93.50 92.10 94.20 97.00 93.80 92.50 94.10 96.90 93.65 92.30
IV-C CycleGAN Enhancement Effects
To improve model performance on night-time images, we employed CycleGAN to convert low-quality night frames to higher-quality day-like images. Using a separate test set of 726 frames that were manually labeled across the three weather classes, we measured the classification performance with and without CycleGAN enhancement.
Figure 4: Overall accuracy comparison across model configurations showing EVA-02+CycleGAN achieving the highest performance at 97.01%, followed by Vision-SigLIP-2+Text-SigLIP-2+CycleGAN+Contrastive at 94.0%, while the baseline Vision Transformer performs poorly at only 55.81%. Models with CycleGAN preprocessing consistently outperform their baseline counterparts.
Figure 5: Class-wise accuracy breakdown across weather conditions demonstrating that while EVA-02+CycleGAN excels in No Precipitation and Rain classes, the Vision-SigLIP-2 models provide more balanced performance across all weather conditions, especially for the challenging Snow class where domain adaptation through CycleGAN and contrastive learning proves particularly beneficial.
Our experimental results, visualized in Figures 4 and 5, demonstrate significant performance improvements when integrating CycleGAN, especially for night-time data. The Vision-SigLIP-2 + text-SigLIP-2 model exhibits substantial gains with CycleGAN integration, improving accuracy by 14 percentage points (from 67.00% to 81.00%), F1 score by 14.25 percentage points (from 66.50% to 80.75%), and precision by 14.50 percentage points (from 66.00% to 80.50%). Similarly, the EVA-02 model shows remarkable enhancement, with accuracy increasing by 19.05 percentage points (from 63.40% to 82.45%), F1 score by 19.56 percentage points (from 62.70% to 82.26%), and precision by 20.00 percentage points (from 62.10% to 82.10%). The significant night-time performance gap observed in the baseline models aligns with findings from previous studies that highlight the challenges of classification under low-light conditions [24, 15].
Interestingly, the Vision Transformer shows minimal improvement with CycleGAN integration, with only a 1.70 percentage point increase in accuracy, no change in F1 score, and a 2.12 percentage point decrease in precision. This further confirms that the Vision Transformer’s limitations with traffic camera imagery extend beyond illumination issues to more fundamental challenges in feature learning from this domain. This finding is consistent with research suggesting that standard ViT models without proper pretraining or data augmentation may struggle with domain-specific tasks on smaller datasets [22, 8].
TABLE III: Impact of CycleGAN enhancement on night-time performance metrics across model architectures. The table shows performance before enhancement (Base), after enhancement (w/ CycleGAN), and the absolute improvement (Diff). EVA-02 and Vision-SigLIP-2 models show substantial gains across all metrics, while Vision Transformer shows minimal or negative changes.
Accuracy (%) Precision (%) F1 Score (%)
Model Base w/ CycleGAN Diff Base w/ CycleGAN Diff Base w/ CycleGAN Diff
EVA-02 63.40 82.45 +19.05 62.10 82.10 +20.00 62.70 82.26 +19.56
Vision-SigLIP-2 + Text-SigLIP-2 67.00 81.00 +14.00 66.00 80.50 +14.50 66.50 80.75 +14.25
Vision Transformer 52.50 54.20 +1.70 63.54 61.42 -2.12 59.96 59.96 0.00
Table III presents a comprehensive view of the night-time performance metrics for all evaluated models. The data clearly demonstrates the effectiveness of our domain adaptation approach for both the EVA-02 and SigLIP-2 models, while highlighting the limitations of the standard Vision Transformer for this application. EVA-02 shows the most dramatic improvements after CycleGAN enhancement, with nearly 20 percentage point gains across all metrics, followed by the Vision-SigLIP-2 models with approximately 14 percentage point improvements. In stark contrast, the Vision Transformer shows negligible or even negative changes, confirming our hypothesis that domain adaptation techniques are most effective when paired with models that have strong feature representation capabilities.
IV-D Qualitative Analysis of Night-to-Day Conversion
Figure 6 provides a visual example of the CycleGAN’s night-to-day conversion process. The original night-time frame (left) shows poor visibility with limited contrast and detail, while the converted day-time image (right) exhibits enhanced illumination and visibility of road features while preserving the original scene’s weather characteristics. The transformation successfully addresses key challenges identified in nighttime traffic imagery [15, 24], including uneven illumination, low contrast, and color distortion, without introducing artificial weather artifacts that could mislead the classifier. This improvement in visual clarity directly translates to better feature extraction by the classification models, as quantitatively demonstrated in our performance analysis.
Figure 6: Example of CycleGAN-based night-to-day conversion showing domain adaptation capabilities: The left image shows the original night-time frame with poor visibility, limited contrast, and color distortion—typical challenges in low-light traffic monitoring. The right image displays the CycleGAN-enhanced version with significantly improved illumination, better contrast, and clearer visualization of critical weather-related features such as road conditions and precipitation patterns. Note how the transformation maintains scene geometry and weather characteristics while enhancing only illumination-dependent features, enabling better downstream classification without introducing misleading artifacts that could affect classification accuracy.
IV-E Computational Efficiency Analysis
Beyond accuracy, computational efficiency is a critical factor for practical deployment in transportation infrastructure systems. Table IV presents a comparison of the computational requirements for the evaluated models. The results highlight a substantial efficiency advantage for our SigLIP-2-based approach, which requires only 40 minutes for training compared to 6 hours for EVA-02—an 89% reduction in training time. Similarly, the inference time on our test set was reduced from approximately 3 minutes for EVA-02 to just 30 seconds for our SigLIP-2-based model, representing an 83% improvement.
TABLE IV: Computational performance comparison across model architectures. Our proposed SigLIP-2 with CycleGAN and contrastive learning framework demonstrates significant efficiency advantages over both EVA-02 and the Vision Transformer models, with substantially reduced training and inference times.
Model Training Time Inference Time
Vision Transformer 6 hours 3 minutes
EVA-02 6 hours 3 minutes
SigLIP-2 + CycleGAN+Constrastive 40 minutes 30 seconds
This dramatic improvement in computational efficiency while maintaining competitive accuracy (94.00% overall accuracy vs. 97.01% for EVA-02) represents a crucial advantage for real-world applications, where deployment on edge devices or resource-constrained systems is often necessary. The reduced computational footprint also translates to lower energy consumption and infrastructure costs, making our approach more sustainable and economically viable for widespread deployment.
V Conclusion and Future Work
In this study, we proposed a robust weather classification framework for low-quality traffic camera imagery, effectively addressing the performance degradation during nighttime conditions by combining domain adaptation through CycleGAN with Transformer-based models. Our results demonstrate that CycleGAN-enhanced domain transformations significantly improve model performance at night, with the Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + contrastive training configuration achieving the best night-time performance (85.90% accuracy) while maintaining high overall accuracy (94.00%). We showed that replacing computationally intensive CLIP with the more efficient SigLIP-2 maintains high accuracy while reducing computational demands—critical for widespread deployment.
The computational efficiency gains are particularly noteworthy, with our approach reducing training time by 89% and inference time by 83% compared to EVA-02, making it substantially more viable for resource-constrained environments. Our approach reduced the performance gap between day and night conditions from 33.81 to 8.90 percentage points in our best model, enabled by our novel weather-preserving loss in the CycleGAN framework that maintains critical visual cues during translation.
Despite these advances, limitations remain: the system struggles with extremely degraded images having near-zero illumination or severe blurring, and our relatively limited dataset constrains the models’ ability to capture rare weather patterns. Future work should explore multi-modal sensing with complementary imaging technologies, explicit temporal modeling through sequence-based architectures, expanded datasets, and edge deployment optimizations. This work addresses transportation safety challenges by providing a cost-effective solution for early, localized weather detection from existing camera infrastructure, potentially reducing weather-related traffic accidents through timely alerts. By addressing the day-night performance gap and reducing computational requirements, our approach offers a viable path toward widespread implementation of camera-based weather monitoring systems for safer transportation networks.
ACKNOWLEDGMENT
The authors would like to express their sincere gratitude to Pelmorex Weather Corporation for funding this research.